# A coupling-based approach to $f$-divergences diagnostics for Markov chain Monte Carlo

### v1

@@ 1.2 Contributions
In this article we propose a novel way to run parallel MCMC chains which gives a consistently weighted approximation of $\pi$ at each time step. These can be interpreted as noisy approximations of the Radon–Nikodym derivative $\mathrm{d}\pi/\mathrm{d}\mu_{t}$. We leverage this construction to introduce upper bounds to the $f$-divergences of the target with respect to the current sample distribution. Contrary to existing diagnostics, our method is fully online and valid from step one, requiring no lag or warm-up. A key consequence stemming from the connection between the $\chi^{2}$-divergence and the effective sample size, is that we can estimate the ‘effective number of active chains’, a highly interpretable and practical diagnostic for practitioners. Our construction is based on Markov kernel coupling techniques arising from the MCMC (Glynn and Rhee, , 2014; Jacob et al., 2020b, ) literature and a weighting structure arising from sequential Monte Carlo literature (see e.g. Chopin and Papaspiliopoulos, , 2020), in particular Dau and Chopin, (2023, Section 4.2) using two interacting copies of the $N$ chains.
The rest of the article is organized as follows. Section 2 explains how any weighted approximation of the target can be used as a diagnostic and highlight the challenge of finding a good one in the context of MCMC. Section 3 details our method, termed weight harmonization, and explains how it can be used both to diagnose the convergence of MCMC and to produce a consistent estimate of expectations under the target for any finite $t$. Section 4 provides consistency and unbiasedness results for the resulting importance weighted MCMC chains. Additionally, under strong mixing assumptions and a coupling assumption akin to uniform ergodicity, we prove that the system of weights converges exponentially fast to uniform weighting of all trajectories, recovering a computational version of exponential ergodicity theorems and validating the use of our method as a diagnostic. Section 5 illustrates the empirical behaviour of our method and show that it returns practical albeit conservative estimators of the effective sample size of the system. Section 6 concludes with a discussion of different advantages and drawbacks of our method. In particular, we identify two likely points of improvement: increasing the interaction between particles by Rao–Blackwellization, and offline correction of the estimators, in a fashion similar to particle smoothing operations (see, e.g., Dau and Chopin, , 2023; Chopin and Papaspiliopoulos, , 2020, Chapter 12).
@@ 2 Importance weighting and Markov chain
A central object in our methodology is the class of weighted approximations of the target distribution. Informally we approximate the target distribution by a system of draws $X^{1},\ldots,X^{N}$ where each particle $X^{i}$ is associated with a weight $W^{i}$ representing its importance. We show how such representations can be obtained in MCMC, how they can produce convergence diagnostics in terms of $f$-divergences, and why advanced methods for setting the weights $W^{i}$ are necessary.
@@ 2.2 Weighted approximations as diagnostics
Let $X^{1},\ldots,X^{N}$ be $N$ (not necessarily independent) identically distributed draws from a distribution $\mu$. We are most interested in the case where $\mu=\mu_{t}=\mu_{0}K^{t}$ is the distribution of a Markov chain state after $t$ iterations of a $\pi$-invariant MCMC kernel $K$ when started at $\mu_{0}$. Suppose further that we have $N$ random weights $W^{1},\ldots,W^{N}$ such that $\sum_{n=1}^{N}W^{n}=1$ and
@@ 2.3 Naive approximation in MCMC and its suboptimality
$\sum_{i=1}^{N}W_{0}^{n}\varphi(X_{t}^{n})\to\int\varphi(x)\pi(\mathrm{d}x)$
@@ 3.1 Couplings of Markov kernels
The key idea of our method is the following: give a set of $N$ chains, we can make them interact pairwise via coupled kernels. When two Markov chains, associated with different weights, eventually collapse to the same state (we say they meet or couple), they can then split even their previous weights and be reallocated to other peer chains. As the simulation progresses, more meeting events will take place between different chains and therefore equalize the weights of all chains present.
@@ 5.1 A fully tractable system
All experiments target the 100-dimensional standard Gaussian, and start from $\mu_{0}\sim\mathcal{N}(101_{100},5I_{100})$, largely outside of stationarity. Confidence intervals are shown as two standard deviations for ten independent realizations of Algorithm 1.
@@
Other $f$-divergences profiles are reported in Appendix J: an interesting point to notice therein is that the Kullback–Leibler divergence bounds exhibit high variance, which is likely a drawback of (3) relying on log-weights for these.
@@ 6.1 Summary
Several properties of the construction have been investigated: first, we (i) have shown consistency at any point in time for increasing number of particles. This is a desirable property that ensures that we can accept the outcome of our algorithm as a trustworthy representation of the target distribution. Second, we (ii) proved that, under assumptions akin to uniform ergodicity, our Radon–Nikodym estimator recovered the expected convergence behaviour: as the number of iteration increases, our weights homogeneize towards a common average value and their variance decreases exponentially fast. Finally, based on a novel bound of $f$-divergences, (iii) we introduced diagnostic criteria for any $f$-divergence of the system of Markov chains generated with respect to its target distribution. This covered in particular the $\chi^{2}$ divergence and total variation distance. To the best of our knowledge, our estimator provides the first computable upper bound to $f$-divergences for Markov chains.
@@ Acknowledgement
We thank Pierre E. Jacob for providing details on the implementation of the Pólya-Gamma coupled sampler. AC acknowledges the financial support provided by UKRI for grant EP/Y014650/1, as part of the ERC Synergy project OCEAN. HD would like to acknowledge support from a Singaporean Ministry of Education Tier 2 grant (MOE-T2EP20123-0010).
@@ Appendix B Proof of Theorem 1
Since $L_{N}\to 0$ in probability, the right-hand side tends to 0 as $N\to\infty$.
@@ Appendix E Proof of Theorem 2
$\operatorname{pr}\left(\hat{\theta}_{t}-\theta_{t-1}\geq\varepsilon\right)\leq\frac{1}{2\varepsilon^{2}N}E[(w_{0}^{1})^{2}K^{t}[\varphi^{2}](X_{0}^{1})].$ (36)

### v2

@@ 1.2 Contributions
In this article we propose a novel way to run parallel MCMC chains which gives a consistently weighted approximation of $\pi$ at each time step. These can be interpreted as noisy approximations of the Radon–Nikodym derivative $\mathrm{d}\pi/\mathrm{d}\mu_{t}$. We leverage this construction to introduce upper bounds to the $f$-divergences of the target with respect to the current sample distribution. Contrary to existing diagnostics, our method is fully online and valid from step one, requiring no lag or warm-up. A key consequence stemming from the connection between the $\chi^{2}$-divergence and the effective sample size, is that we can estimate the ‘effective number of active chains’, a highly interpretable and practical diagnostic for practitioners. Our construction is based on Markov kernel coupling techniques arising from the MCMC (Glynn and Rhee, , 2014; Jacob et al., 2020b, ) literature and a weighting structure arising from sequential Monte Carlo literature (see e.g. Chopin and Papaspiliopoulos, , 2020), in particular Dau and Chopin, (2023, Section 4.2) using two interacting copies of $N$ independent chains.
The rest of the article is organized as follows. Section 2 explains how any weighted approximation of the target can be used as a diagnostic and highlights the challenge of finding a good one in the context of MCMC. Section 3 details our method, termed weight harmonization, and explains how it can be used both to diagnose the convergence of MCMC and to produce a consistent estimate of expectations under the target for any finite $t$. Section 4 provides consistency and unbiasedness results for the resulting importance weighted MCMC chains. Additionally, under strong mixing assumptions and a coupling assumption akin to uniform ergodicity, we prove that the system of weights converges exponentially fast to uniform weighting of all trajectories, recovering a computational version of exponential ergodicity theorems and validating the use of our method as a diagnostic. Section 5 illustrates the empirical behaviour of our method and shows that it returns practical albeit conservative estimators of the effective sample size of the system. Section 6 concludes with a discussion of different advantages and drawbacks of our method. In particular, we identify two likely points of improvement: increasing the interaction between particles by Rao–Blackwellization, and offline correction of the estimators, in a fashion similar to particle smoothing operations (see, e.g., Dau and Chopin, , 2023; Chopin and Papaspiliopoulos, , 2020, Chapter 12).
@@ 2.2 Weighted approximations as diagnostics
A central object in our methodology is the class of weighted approximations of the target distribution. Let $X^{1},\ldots,X^{N}$ be $N$ (not necessarily independent) identically distributed draws from a distribution $\mu$. We are most interested in the case where $\mu=\mu_{t}=\mu_{0}K^{t}$ is the distribution of a Markov chain state after $t$ iterations of a $\pi$-invariant MCMC kernel $K$ when started at $\mu_{0}$. Suppose further that we have $N$ random weights $W^{1},\ldots,W^{N}$ such that $\sum_{n=1}^{N}W^{n}=1$ and
@@ 2.3 Naive approximation in MCMC and its suboptimality
$\sum_{n=1}^{N}W_{0}^{n}\varphi(X_{t}^{n})\to\int\varphi(x)\pi(\mathrm{d}x)$
@@ 3.1 Couplings of Markov kernels
The key idea of our method is the following: given a set of $N$ chains, we can make them interact pairwise via coupled kernels. When two Markov chains, associated with different weights, eventually collapse to the same state (we say they meet or couple), they can then split even their previous weights and be reallocated to other peer chains. As the simulation progresses, more meeting events will take place between different chains and therefore equalize the weights of all chains present.
@@ 5.1 A fully tractable system
All experiments target the 100-dimensional standard Gaussian, and start from $\mu_{0}\sim\mathcal{N}(10\cdot 1_{100},5\cdot I_{100})$, largely outside of stationarity. Confidence intervals are shown as two standard deviations for ten independent realizations of Algorithm 1.
@@
Other $f$-divergences profiles are reported in Appendix J: an interesting point to notice therein is that the reverse Kullback–Leibler divergence bounds was largely too conservative and exhibited high variance, which is likely an empirical counterpart to Corollaries 3 and 4 not applying. As a consequence we did not report it there.
@@ 6.1 Summary
Several properties of the construction have been investigated: first, we (i) have shown consistency at any point in time for increasing number of particles. This is a desirable property that ensures that we can accept the outcome of our algorithm as a trustworthy representation of the target distribution. Second, we (ii) proved that, under assumptions akin to uniform ergodicity, our Radon–Nikodym estimator recovered the expected convergence behaviour: as the number of iteration increases, our weights homogeneize towards a common average value and their variance decreases exponentially fast. Finally, based on a novel bound of $f$-divergences, we (iii) introduced diagnostic criteria for any $f$-divergence of the system of Markov chains generated with respect to its target distribution. This covered in particular the $\chi^{2}$ divergence and total variation distance. To the best of our knowledge, our estimator provides the first computable upper bound to $f$-divergences for Markov chains.
@@ Acknowledgement
We thank Pierre E. Jacob for providing details on the implementation of the Pólya-Gamma coupled sampler and Yvann Le Fay for saving a factor 2 from being forgotten. AC acknowledges the financial support provided by UKRI for grant EP/Y014650/1, as part of the ERC Synergy project OCEAN. HD would like to acknowledge support from a Singaporean Ministry of Education Tier 2 grant (MOE-T2EP20123-0010).
@@ Appendix B Proof of Theorem 1
Since $L_{N}\to 0$ in probability, the right-hand side tends to 0 as $N\to\infty$. ∎
@@ Appendix E Proof of Theorem 2
$\operatorname{pr}\left(\hat{\theta}_{t}-\theta_{t-1}\geq\varepsilon\right)\leq\frac{1}{\varepsilon^{2}N}E[(w_{0}^{1})^{2}K^{t}[\varphi^{2}](X_{0}^{1})].$ (36)